A person lives on the third floor of a building and he runs up the stairs three times a day. Each floor has 24 steps and height of each step is 15 cm. a) If the person weight is 65 kg calculate how much energy he uses each day. b) If he runs from first floor to the third floor in 20 seconds, calculate the person power output
Answer #1

A)

Energy uses per day = 3 * mg h = 3 * 65* 9.8* 0.15* 24

E = 6879.6 J

B)

P = E/t = 6879.6 / 20 = 343.98 W

Comment in case any doubt please rate my answer....